OpenStudy (anonymous):

write an equation of a circle with the given center and radius. center (2, 4) and radius 5 (x – 2)2 + (y + 4)2 = 5 (x + 2)2 + (y – 4)2 = 5 (x + 2)2 + (y – 4)2 = 25 (x – 2)2 + (y – 4)2 = 25

OpenStudy (anonymous):

\[(x-x _{1}^{})^2+ (y-y _{1}^{})^2 = r^2\]

OpenStudy (anonymous):

so x1 =2 and y1 =4 and r=5

OpenStudy (anonymous):

just put the values

OpenStudy (anonymous):

D....thank you!

OpenStudy (anonymous):

Help me with another one please??! :) find an equation of a parabola with a vertex at the origin and directrix y = –1.5.

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

x =1/6 –y2 x =1/6 y2 y =1/6 –x2 y =1/6 x2

OpenStudy (anonymous):

where p is the directrix

OpenStudy (anonymous):

\[x^2=4py\]

OpenStudy (anonymous):

x=1/6-y^2

OpenStudy (anonymous):

no its y=-x^2/6

OpenStudy (anonymous):

\[x^2= 4py .....x^2=-6y .....\] divide by -6 \[-x^2/6 =y\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yea thx

OpenStudy (anonymous):

originally it would be x=1/6y^2?

OpenStudy (anonymous):

wait wait

OpenStudy (anonymous):

y =1/6 –x2 is this option \[y=\frac{ 1 }{ 6 }(-x)^2 \] then it is correct

OpenStudy (anonymous):

no something is wrong write the options with LateX

OpenStudy (anonymous):

write them using equation